Equal amounts are invested in each of the three accounts paying 7%, 8% and 10.5% annually. If one year's combined interest income is $1249.50, how much is invested in each amount?
Answer by geetha_rama(94)

Let x be the amount invested in each of the three accounts
Interest earned in account paying 7% interest I1 = 7x/100
Interest earned in account paying 8% interest I2 = 8x/100
Interest earned in account paying 10.5% interest I3 = 10.5x/100
Total Interest = 1249.50
I1 + I2 +I3 = 1249.50
7x/100 + 8x/100 + 10.5/100 = 1249.50
=> 7x + 8x + 10.5x = 124950
=>25.5x = 124950
x = 4900
Amount investeed in each of the account is $4900